OpenStudy (anonymous):

How would you simplify a radical if it is just a varible? The Square root of p with an exponent of 15

OpenStudy (anonymous):



OpenStudy (anonymous):

is this the qustion????

OpenStudy (anonymous):

its more like \[\sqrt{p}15\]

OpenStudy (aravindg):

u can take \[\sqrt{p}=p^{\dfrac{1}{2}}\]

OpenStudy (anonymous):